A banquet hall offers two types of tables for​ rent: 6-person rectangular tables at a cost of $22 each and​ 10-person round tables at a cost of $53 each. Kelly would like to rent the hall for a wedding banquet and needs tables for 240 people. The room can have a maximum of 39 tables and the hall only has 15 rectangular tables available. How many of each type of table should be rented to minimize cost and what is the minimum​ cost?
Answer #1

Answer:

Cost per person
rectangular = 22/6 =5.5
Round = 53/10 = 5.3
15 maximum rectangular tables
15 x 6 = 90 people
240-90 = 150

Round Tables
150/10 = 15
240 people

90 people with 6 people at each 15 rectangular table at $22 each = 15 x 22 = 330
150 people with 10 people at each of 15 round tables at $53 each = 15 x 53 = 795
Total cost = 330+795 = 1125

The minimum cost $1125, is achieved with 15 6-person rectangular table and 15 10-person round table.